### v1

# Gauss Principle in Incompressible Flow: A Unified Variational Perspective on Pressure and Projection

Abstract
Following recent work (Gonzalez & Taha 2022; Peters & Ormiston 2025), this manuscript clarifies what the Gauss–Appell principle determines in incompressible, inviscid flow and how it connects to classical projection methods. At a fixed time, freezing the velocity and varying only the material acceleration leads to minimization of a quadratic subject to acceleration-level constraints. First-order conditions yield a Poisson–Neumann problem for a reaction pressure whose gradient removes the non-solenoidal and wall-normal content of the provisional residual, i.e. the Leray–Hodge projection. Thus, Gauss–Appell enforces the instantaneous kinematic constraints and recovers Euler at the instant. In steady flows, this principle cannot select circulation or stagnation points because these are properties of the velocity state, not the instantaneous acceleration correction. The principle only determines the reaction pressure for an already-specified velocity field. The impressed/reaction pressure decomposition can be supplemented with orthogonality conditions, allowing physical inputs (e.g. circulation, freestream) to be embedded while maintaining exact constraint enforcement. This variational viewpoint provides a diagnostic for computational incompatibility: In a practical computation, spikes in projection effort may signal problematic boundary conditions or under-resolution, and clarifies how pressure instantaneously maintains solenoidality without creating or destroying vorticity. The goal of this note is simply to lend more clarity to the application of the Gauss principle, and to connect it concretely to well known concepts including potential flow theory, projection algorithms, and recent variational approaches.
keywords:
Variational theory; Principle of least constraint; Gauss principle
The application of variational principles to fluid mechanics has a rich history spanning over 150 years, beginning with Helmholtz’s vortex theorems (1858) and Kelvin’s circulation theorem (1869). Classical approaches typically employed Hamilton’s principle, extremizing action integrals over time to derive the equations of motion through calculus of variations (Holm et al., 1998; Truesdell & Noll, 2004; Marchioro & Pulvirenti, 2012). These time-integral formulations, while mathematically elegant, vary entire space-time paths and thus provide limited insight into the instantaneous mechanics of constraint enforcement in incompressible flows.  (Holm et al., 1998; Truesdell & Noll, 2004; Marchioro & Pulvirenti, 2012).
A fundamentally different variational approach emerged from Gauss’s principle of least constraint (Gauss, 1829), originally developed for constrained mechanical systems. Unlike Hamilton’s principle which varies paths over time, Gauss’s principle operates at a fixed instant, minimizing a quadratic form in accelerations subject to acceleration-level constraints. This instantaneous character makes it particularly suited to understanding how pressure maintains incompressibility moment by moment. Despite its potential, the principle remained largely unexplored in fluid mechanics for over a century. Recent works (Gonzalez & Taha, 2022; Taha & Shorbagy, 2025; Peters & Ormiston, 2025) have brought back a variational lens on fluid flows, particularly revisiting the Gauss principle in the context of the airfoil lift problem.
Despite the renewed interest, there remains confusion about what a Gauss/Appell least‑constraint statement actually determines in fluids, and how it can be utilized. We provide a compact variational interpretation of the pressure Poisson equation, clarify how to embed physical inputs via an ‘impressed’ subspace, and give a corollary that places the projection method in numerical hydrodynamics (Chorin, 1968) under Gauss/Appell.
Gonzalez & Taha (2022) minimize a steady reduced functional on a kinematically admissible manifold (e.g., a potential‑flow family parameterized by circulation), a legitimate model‑reduction device but not the instantaneous Gauss principle. Taha & Shorbagy (2025) recast this as a “minimum pressure‑gradient” criterion as a selection mechanism for weak Euler solutions—again a steady or quasi‑steady choice layered on top of the equations. Peters & Ormiston (2025) correctly argue that least constraint, by itself, cannot resolve classical indeterminacies of steady exterior potential flow.
This work reconciles these viewpoints by showing precisely where Gauss/Appell ends (fixed‑time projection) and where additional modeling/selection begins (steady‑state choice of manifold or boundary data). The Gauss principle is a fixed-time variational statement: one freezes the velocity field and varies the material acceleration. Minimizing a positive‑definite “Appellian” in the admissible accelerations yields first order optimality conditions that are exactly the Leray–Hodge projection of the residual onto gradient fields; the Lagrange multiplier is a reaction pressure obtained from a Poisson–Neumann problem whose gradient removes the non‑solenoidal part of the residual. In steady flow, this principle does not select among kinematically admissible steady states; it only determines the pressure compatible with the given velocity field. These points are well known from projection methods but are often obscured in recent discussions; we make the equivalence and its scope precise.
Further, the non-uniqueness of the impressed vs. constraint split is clarified, and the decomposition of Peters & Ormiston (2025) is substantiated via an energy-orthogonality condition that selects a unique decomposition relative to a chosen impressed subspace, and by connecting this result to classical projection. It is shown however, that while physically intuitive choices can be made for the impressed subspace, the choice itself is arbitrary.
Unlike the principle of least action, Gauss/Appell is a least‑deviation principle in the accelerations at a given instant. In this sense, it is clarified that Gauss/Appell provides a unifying variational interpretation of the pressure Poisson solve—useful analytically and practically, particularly when one wishes to embed physical inputs (e.g., circulation or far‑field content) through an impressed subspace while keeping the constraints exact. We present rigorous derivations in both Eulerian and Lagrangian, and provide insight into the physical implications of the Gauss principle.
1 Setting, notations
Let $\Omega$ be the exterior domain of a smooth body, and $\partial\Omega$ be the boundaries. The density $\rho$ is constant. $u(\cdot,t)$ is the velocity field. The Gauss’ principle is applied at a fixed time. One freezes the state $u(\cdot,t)$ and chooses an admissible instantaneous update in acceleration. For fluids, the material acceleration is $a\triangleq\partial_{t}u+(u\!\cdot\!\nabla)u.$ At a fixed time, the convective part $(u\!\cdot\!\nabla)u$ is a known function of space (because $u(\cdot,t)$ is frozen). Hence a perturbation of acceleration comes entirely from perturbing $u_{t}$: $\delta a=\delta u_{t}$. This is the fluid analogue of Gauss in constrained multibody dynamics (Udwadia & Kalaba, 1992), where one minimizes a positive-definite quadratic form in accelerations at a fixed configuration and velocity, subject to acceleration-level constraints; here the Eulerian fixed-time data are $(u,\nabla\!\cdot u,u\!\cdot n)$. Therefore, the variational variable is $u_{t}(\cdot,t)$, not a time-slab history, and not a configuration displacement as in stationary action.
For an airfoil, with far-field velocity $U_{\infty}$ and slip wall:
$\nabla\!\cdot u=0\quad\text{in }\Omega,\qquad u\!\cdot n=0\quad\text{on }\partial\Omega,\qquad u\to U_{\infty}\text{ as }|x|\to\infty.$
At the fixed instant $t$, these are given. Differentiating the two kinematic constraints in time gives the acceleration-level constraints on $u_{t}$:
$\nabla\!\cdot u_{t}=0\quad\text{in }\Omega,\qquad u_{t}\!\cdot n=0\quad\text{on }\partial\Omega.$ (1)
Below, the Eulerian form of the derivation is detailed. The Lagrangian form can be found in the appendix.
1.1 Impressed vs reaction pressure and the Appellian
As in Peters & Ormiston (2025), we split the total pressure
$p\triangleq P_{F}+P_{R},$
where
  • $P_{F}$ (impressed) collects the gradient content we treat as “given” at the instant and that does not necessarily enforce the kinematic constraints.
  • $P_{R}$ (constraint) is the Lagrange–multiplier that enforces the acceleration–level constraints completely and uniquely. Its gradient $\nabla P_{R}$ is the constraint force per unit volume.
Define the frozen field
$C\triangleq(u\!\cdot\!\nabla)u+\frac{1}{\rho}\,\nabla P_{F},$ (2)
which is entirely known once $u(\cdot,t)$ and $P_{F}(\cdot,t)$ are specified. Gauss/Appell at the instant minimizes the Appellian
[…]
2 First order stationarity
Construct the Lagrangian with multipliers $\lambda$ (volume) and $\mu$ (boundary):
$\mathcal{L}[u_{t},\lambda,\mu]\triangleq\frac{1}{2}\int_{\Omega}\rho\,|u_{t}+C|^{2}\,dV\;-\;\int_{\Omega}\lambda\,(\nabla\!\cdot u_{t})\,dV\;-\;\int_{\partial\Omega}\mu\,(u_{t}\!\cdot n)\,dS.$
Variations yield
$\delta_{\lambda}:\ \nabla\!\cdot u_{t}=0,\qquad\delta_{\mu}:\ u_{t}\!\cdot n=0,$
and
$\delta_{u_{t}}\mathcal{L}$ $=\int_{\Omega}\big(\rho(u_{t}+C)+\nabla\lambda\big)\cdot\delta u_{t}\,dV\;+\;\int_{\partial\Omega}(\lambda-\mu)\,\delta u_{t}\!\cdot n\,dS.$
Because $\delta u_{t}$ is arbitrary in the interior and its normal component is arbitrary on $\partial\Omega$ (before enforcing the constraint), stationarity gives
$\rho(u_{t}+C)+\nabla\lambda=0\quad\text{in }\Omega,\qquad\mu=\lambda\quad\text{on }\partial\Omega,$ (4)
together with (1). This derivation is aligned with  Peters & Ormiston (2025), though we clarify the frozen field C.
Now, define $P_{R}\triangleq\lambda$. From (4), $u_{t}=-(C+\nabla P_{R}/\rho)$; impose (1):
$\nabla\!\cdot u_{t}=0\ \Rightarrow\ \nabla^{2}P_{R}=-\,\rho\,\nabla\!\cdot C\quad\text{in }\Omega,$
and
$u_{t}\!\cdot n=0\ \Rightarrow\ \partial_{n}P_{R}=-\,\rho\,C\!\cdot n\quad\text{on }\partial\Omega.$
Thus the first step is the scalar Poisson–Neumann boundary value problem
$\boxed{\;\nabla^{2}P_{R}=-\,\rho\,\nabla\!\cdot C\ \text{ in }\Omega,\qquad\partial_{n}P_{R}=-\,\rho\,C\!\cdot n\ \text{ on }\partial\Omega}$ (5)
Compatibility holds identically: $\int_{\Omega}\nabla^{2}P_{R}\,dV=\int_{\partial\Omega}\partial_{n}P_{R}\,dS$ because $\int_{\Omega}\nabla\!\cdot C\,dV=\int_{\partial\Omega}C\!\cdot n\,dS.$ With (5) solved, we have $u_{t}^{\star}=-(C+\nabla P_{R}/\rho),\qquad p=P_{F}+P_{R}.$ Since $a=u_{t}+(u\!\cdot\!\nabla)u$, (4) is equivalent to
$\rho a=-\nabla p,$
i.e. Euler at the instant.
Once optimized, the minimized Appellian is purely “projection effort”
$\mathcal{A}_{\star}=\frac{1}{2}\int_{\Omega}\frac{1}{\rho}\,|\nabla P_{R}|^{2}\,dV.$
This is precisely the standard projection method logic in continuous form: solve a Poisson–Neumann problem for pressure and project the residual onto gradients. It is notable that the Leray projector (see, for instance, Temam (2024)) maps any field to the divergence-free, wall-tangent subspace. Writing $R\triangleq\rho\ a+\nabla P_{F}=\rho(u_{t}+C)+\nabla P_{F},$ the KKT conditions $\rho(u_{t}^{\star}+C)=-\nabla P_{R}$ says $R=-\nabla P_{R}$: the solenoidal, wall-tangent part of $R$ vanishes. Thus $P_{R}$ realizes the Leray projection of $R$ onto gradients, and 5 is the strong-form projector.
3 Physical implications and insight
At a fixed instant, the Gauss–Appell step has a concrete physical meaning: it chooses the smallest admissible change to the material acceleration that restores the instantaneous momentum balance while preserving incompressibility and wall non-penetration. The only agency it has is a reaction field, $\nabla P_{R}/\rho$, whose role is to remove whatever divergence and normal flux the provisional residual carries. The pressure is a reaction that instantaneously enforces the kinematics; it does no work on divergence-free, wall-tangent motions. This point is correctly argued by  Gonzalez & Taha (2022). After elimination, the minimized Appellian $\mathcal{A}^{\ast}$ becomes a quantitative measure of the projection effort, i.e. how hard the constraint must “push” to keep the instantaneous update compatible with incompressibility and wall tangency.
3.1 Steady flow
It is emphasized that in a steady flow, the Gauss principle of least constraint finds the gradient one must subtract from the “residual” $\rho(u\cdot\nabla)u+\nabla P_{F}$ so that the momentum equation holds. There is no update to $u$ (the minimizer gives $u_{t}^{\ast}=0$) and no selection among multiple steady Euler fields that satisfy the kinematics. In other words, no selection of extra physics occurs.
Gonzalez & Taha (2022) restrict the velocity field to a steady potential-flow family $u(\cdot;\Gamma)\triangleq u_{0}+\Gamma u_{1}$ whose members satisfy continuity and wall tangency, and then define
$S(\Gamma)\triangleq\frac{1}{2}\int_{\Omega}\rho\,\big|\,u\!\cdot\!\nabla u\,\big|^{2}\,dV,$
seeking $\Gamma^{\star}=\arg\min_{\Gamma}S(\Gamma)$. This is a legitimate reduced-model optimization on a one-parameter manifold of kinematically admissible steady flows and can, in practice, pick a $\Gamma$ consistent with certain steady configurations. It can be used as a closure condition. However, it is not the instantaneous Gauss/Appell principle; it presupposes steadiness and a particular parameterization. In the full Gauss setting, the state $u$ is frozen, the variable is $u_{t}$, and the outcome is the constraint projection $P_{R}$2), not the selection of $\Gamma$. In the steady case, for every $\Gamma$ there exists a pressure that makes $(u(\cdot;\Gamma),p(\cdot;\Gamma))$ a steady Euler solution. Applying Gauss at the instant then yields $u_{t}^{\ast}=0$ for every $\Gamma$. Therefore, the principle does not (and cannot) decide which $\Gamma$ nature takes. In fact, in the steady flow case, the application of the Gauss principle does not yield any new information other than the fact that the reaction pressure can be determined.
Taha & Shorbagy (2025) propose the ‘Principle of Minimum pressure gradient’, and offer useful reduced‑manifold selections that align with classical expectations (Kutta condition on nearly sharp airfoils, Glauert’s rapidly rotating cylinder, empirical separation angles). An important distinction is required to interpret scope: Gauss/Appell provides the fixed‑time projection; selection over steady states requires additional modeling choices.
In the unsteady case, since $u_{t}$ (hence $a$) is varied with $u$ held fixed, acceleration-level constraints are enforced, resulting in the reaction pressure via the Poisson–Neumann problem (5). This exactly recovers Euler at the instant and clarifies the roles of impressed vs constraint pressure. As argued by  Peters & Ormiston (2025), this also explains why least constraint cannot, by itself, resolve the classical indeterminacy of steady exterior potential flow. That indeterminacy lives in the state manifold, not in the instantaneous constrained acceleration update.
3.2 Impressed/constraint split is arbitrary
Let $\psi$ be a smooth field (e.g. $\psi=\Gamma\,\psi_{s}$ representing a circulation mode). Define a new split $\widetilde{P}_{F}\triangleq P_{F}+\psi,\quad\widetilde{P}_{R}\triangleq P_{R}-\psi.$ Then
$\widetilde{C}=(u\!\cdot\!\nabla)u+\frac{1}{\rho}\nabla\widetilde{P}_{F}=C+\frac{1}{\rho}\nabla\psi,$
and the first order optimality condition becomes
$\rho\,(u_{t}^{\star}+\widetilde{C})+\nabla\widetilde{P}_{R}=\rho\,(u_{t}^{\star}+C)+\nabla P_{R}=0,$
so the optimizer $u_{t}^{\star}$ is invariant. Likewise $p=\widetilde{P}_{F}+\widetilde{P}_{R}$ is unchanged. Hence putting the circulation parameter $\Gamma$ “inside $P_{R}$” or “inside $P_{F}$” produces the same physical $(u_{t}^{\star},p)$; the split has a one‑parameter family of representatives. This is why $P_{R}$ is not unique a priori.
Similarly, the Poisson equation changes to $\nabla^{2}\widetilde{P}_{R}=-\rho\,\nabla\!\cdot C-\nabla^{2}\psi,\ \partial_{n}\widetilde{P}_{R}=-\rho\,C\!\cdot n-\partial_{n}\psi$. Therefore $\widetilde{P}_{R}=P_{R}-\psi+\text{const}$, exactly as above. So the Poisson step leaves $(u_{t}^{\star},p)$ invariant. Thus, in incompressible Euler at a fixed time, any decomposition $p=P_{F}+P_{R}$ leaves momentum balance unchanged. This simple observation allows one to assign an impressed component $P_{F}$ and let the constraint $P_{R}$ enforce the acceleration‑level constraints, with no change to the physical solution. Once $P_{R}$ is solved, the optimal update is $u_{t}^{\star}=-(C+\nabla P_{R}/\rho)$, and the instantaneous Euler balance $\rho(\partial_{t}u+(u\!\cdot\!\nabla)u)=-\nabla(P_{F}+P_{R})$ is recovered.
3.3 Uniqueness of Dirichlet Orthogonality
Peters & Ormiston (2025) address the above arbitrariness by proposing an intuitive orthogonal decomposition
$\int_{\Omega}\nabla P_{R}\cdot\nabla P_{F}\,dV=0$ (6)
In Appendix A, it is shown that given any provisional solution $(\widehat{P}_{F},\widehat{P}_{R})$, the unique orthogonal pair is obtained by a correction:
$P_{R}=\widehat{P}_{R}-\sum_{j}\alpha_{j}\psi_{j},\qquad P_{F}=\widehat{P}_{F}+\sum_{j}\alpha_{j}\psi_{j}.$
The coefficients $\alpha_{j}$ are the orthogonal projection of $\widehat{P}_{R}$ onto the impressed subspace $\mathcal{S}_{\text{imp}}=\text{span}\{\psi_{1},\ldots,\psi_{m}\}$. We are subtracting this projection from $P_{R}$ (making it orthogonal to $\mathcal{S}_{\text{imp}}$) and adding it to $P_{F}$ (so the total pressure remains unchanged).
In other words, one can choose
  1. 1.
    What $\mathcal{S}_{\text{imp}}=\text{span}\{\psi_{1},\ldots,\psi_{m}\}$ is — which pressure modes one may want to impress (e.g., circulation patterns)
  2. 2.
    How much of each mode to impress, typically by specifying boundary conditions or forcing terms
The Dirichlet orthogonality condition then ensures consistent bookkeeping when decomposing the resulting total pressure $p$ (which emerges from the full flow physics). $\mathcal{S}_{\text{imp}}$ is the “signature” of the forcing mechanism. Dirichlet orthogonality thus labels as forcing ($P_{F}$) everything that has this signature; and labels as constraint ($P_{R}$) everything orthogonal to it (in the Dirichlet sense). We further note that the above is exactly analogous to the Helmholtz–Leray projection operator $\mathbb{P}$ operating on any given velocity field $w$:
$\mathbb{P}w\;=\;w-\nabla q,\qquad\nabla^{2}q=\nabla\!\cdot w\ \ \text{in }\Omega,\qquad\partial_{n}q=w\!\cdot\!n\ \ \text{on }\partial\Omega.$
Equivalently, $\mathbb{P}w$ is the unique vector that satisfies orthogonality $\int_{\Omega}\mathbb{P}w\cdot\nabla\phi\,\mathrm{d}x=0$. This is a well-known result (See for instance Sohr (2012)), but the connection to the impressed/constraint split is new. Appendix B clarifies another point relevant to the projection.
3.4 Physical choice of impressed subspace
For the exterior airfoil problem the physical implications are immediate. If $S_{\rm imp}$ contains the potential-flow circulation mode, taking $P_{F}=\Gamma\,\psi_{s}$ neatly encodes the circulation input while leaving $P_{R}$ to enforce the constraints for the actual frozen velocity. However, we now show that even with Dirichlet orthogonality, the choice of impressed subspace is highly arbitrary.
  1. 1.
    Circulation basis . Choose $\mathcal{S}_{\!\rm imp}=\mathrm{span}\{\psi_{s}\}$ so that $\nabla\psi_{s}$ represents the potential–flow circulation mode (the harmonic field tangent to the wall with unit loop integral). Then $P_{F}=\Gamma\,\psi_{s}$ carries precisely the circulation degree of freedom emphasized in Gonzalez & Taha (2022), and $P_{R}$ enforces the constraints. This matches the “circulation‑aware” split described in  Peters & Ormiston (2025).
  2. 2.
    Freestream split. It is also legitimate to designate the freestream direction as impressed. A minimal way is to include the linear potential $\phi_{\infty}(x)=U_{\infty}\!\cdot x$ among the generators and let $\psi_{\infty}:=\rho\,\phi_{\infty}$. Then $\nabla P_{F}$ may contain the constant vector $\rho U_{\infty}$. This merely tilts the “target” in $C=(u\!\cdot\!\nabla)u+\nabla P_{F}/\rho$; the Poisson step automatically compensates so that $u_{t}$ remains divergence‑free and wall‑tangent. There is no requirement that this impressed piece satisfy the wall condition—it need not, because only $u_{t}$ is constrained.
  3. 3.
    Full potential reference. One may also place the entire steady Bernoulli gradient for a chosen reference potential flow into $P_{F}$ (e.g. the zero‑circulation solution $u_{0}$ or $u_{0}+\Gamma u_{1}$). In steady irrotational settings, $(u\!\cdot\!\nabla)u=\nabla(\tfrac{1}{2}|u|^{2})$; taking $P_{F}=\rho\,\tfrac{1}{2}|u_{\rm ref}|^{2}$ pushes that Bernoulli content into the impressed side and leaves $P_{R}$ to enforce the constraints for the actual frozen $u$.
The latter two results are particularly important given the criticism of  Peters & Ormiston (2025) of  Gonzalez & Taha (2022)’s derivation as incorrect because they ignored impressed pressure. While in some cases (e.g. accounting for body forces due to gravity), impressed forces can be clearly delineated, it is indeed not incorrect to consider the entire field as non-impressed in the typical airfoil scenario. In other words, even if one impresses the uniform stream, or even a full steady potential reference, the same mechanism applies: the Gauss step removes only what is needed to satisfy the constraints and leaves the impressed content intact. In all cases the reaction is irrotational and hence cannot create or destroy vorticity; it instantaneously redistributes pressure to maintain solenoidality and wall compatibility while vorticity evolves through advection (and any impressed physics).
Finally, this viewpoint unifies interpretation and practice. The standard predict–Poisson–correct algorithm is precisely the Gauss–Appell step on a grid, and $\mathcal{A}^{\star}$ furnishes a diagnostic that can be monitored in computation: spikes signal incompatible boundary data, under-resolved features, or an unhelpful choice of impressed content. In reduced-order or data-assisted settings, the choice of $S_{\rm imp}$ allows trusted physics to be injected while the same projection guarantees exact enforcement of incompressibility and wall conditions.
Acknowledgments
Thanks to Robert Ormiston, Haithem Taha, Dave Peters and Philip Roe for discussions, and for inspiring this work.
Appendix A: Uniqueness of Dirichlet orthogonality
Substitute the correction 3.3 into the orthogonality condition 6:
$0=\int_{\Omega}\nabla P_{R}\cdot\nabla\psi_{i}\,dV=\int_{\Omega}\nabla\left(\widehat{P}_{R}-\sum_{j=1}^{m}\alpha_{j}\psi_{j}\right)\cdot\nabla\psi_{i}\,dV$ (7)
$\int_{\Omega}\nabla\widehat{P}_{R}\cdot\nabla\psi_{i}\,dV-\sum_{j=1}^{m}\alpha_{j}\int_{\Omega}\nabla\psi_{j}\cdot\nabla\psi_{i}\,dV$ (8)
Rearranging gives the $m\times m$ linear system:
$\sum_{j=1}^{m}\alpha_{j}\underbrace{\int_{\Omega}\nabla\psi_{j}\cdot\nabla\psi_{i}\,dV}_{G_{ij}}=\underbrace{\int_{\Omega}\nabla\widehat{P}_{R}\cdot\nabla\psi_{i}\,dV}_{b_{i}},\quad i=1,\ldots,m$ (9)
In matrix form: $G\bm{\alpha}=\mathbf{b}$, where:
  • $G_{ij}=\int_{\Omega}\nabla\psi_{j}\cdot\nabla\psi_{i}\,dV$ is the Gram matrix in the Dirichlet inner product $\langle p,q\rangle\triangleq\int_{\Omega}\nabla p\cdot\nabla q\,dV$.
  • $b_{i}=\int_{\Omega}\nabla\widehat{P}_{R}\cdot\nabla\psi_{i}\,dV$ projects the provisional reaction pressure onto each basis function
Once $\bm{\alpha}$ is computed by solving $G\bm{\alpha}=\mathbf{b}$, the unique Dirichlet-orthogonal decomposition is:
$P_{R}=\widehat{P}_{R}-\sum_{j=1}^{m}\alpha_{j}\psi_{j},\qquad P_{F}=\widehat{P}_{F}+\sum_{j=1}^{m}\alpha_{j}\psi_{j}$ (10)
Appendix B: Rank and Projection
Peters & Ormiston (2025) cast multipliers via a finite‑dimensional Lagrange‑multiplier theorem, tying the number of independent multipliers to the rank of a constraint Jacobian. That language is useful but does not transfer cleanly to the PDE setting of incompressible flow, where the constraint is an operator ($\nabla\!\cdot$) and the multiplier $P_{R}$ is an entire field. What replaces “rank” is the orthogonal projection onto the divergence‑free, wall‑tangent subspace (Leray projector), with the multiplier recovered as the gradient part in the Helmholtz–Hodge decomposition; see the Euler/Appellian equivalence proof in Gonzalez & Taha (2022). In other words, pressure acts as the constraint that removes the non‑solenoidal content of $\rho(u_{t}+C)$, and the correct mathematical object is a projection, and not a finite matrix rank.
Appendix C: Lagrangian form
In contrast to prior literature, we additionally present the derivation in Lagrangian form, as it is more intuitive and applies more naturally to the steady state treatment. Let $\Omega_{0}$ be a reference domain with boundary $\partial\Omega_{0}$. Consider instantaenous particle locations $X(\xi,\eta,t)$, where $\xi,\eta$ are reference coordinates. At the fixed time $t_{0}$, abbreviate
$X(\xi,\eta)\triangleq X(\xi,\eta,t_{0}),\qquad\dot{X}(\xi,\eta)\triangleq\partial_{t}X(\xi,\eta,t_{0}),\qquad a(\xi,\eta)\triangleq\ddot{X}(\xi,\eta,t_{0})\in\mathbb{R}^{2}.$
Kinematic relations:
$J=X_{\xi}Y_{\eta}-X_{\eta}Y_{\xi},\qquad\widetilde{J}=\begin{pmatrix}Y_{\eta}&-\,Y_{\xi}\\ -\,X_{\eta}&X_{\xi}\end{pmatrix},\qquad\partial_{\xi}\widetilde{J}_{i\xi}+\partial_{\eta}\widetilde{J}_{i\eta}=0\ (i=1,2).$ (11)
Time derivatives:
$\dot{J}=\widetilde{J}_{ij}\,\partial_{x_{0}j}\dot{X}_{i},\qquad\ddot{J}=\widetilde{J}_{ij}\,\partial_{x_{0}j}a_{i}+\dot{\widetilde{J}}_{ij}\,\partial_{x_{0}j}\dot{X}_{i},\qquad\dot{\widetilde{J}}=\begin{pmatrix}\dot{Y}_{\eta}&-\,\dot{Y}_{\xi}\\ -\,\dot{X}_{\eta}&\dot{X}_{\xi}\end{pmatrix}$. (12)
Feasibility assumptions at $t_{0}$ (hold before the Gauss step):
  • Incompressibility: $J(\xi,\eta)=1$ in $\Omega_{0}$.
  • Velocity compatibility: $\dot{J}(\xi,\eta)=\widetilde{J}_{ij}\,\partial_{x_{0}j}\dot{X}_{i}=0$ in $\Omega_{0}$.
  • Wall non–penetration: $\dot{X}\!\cdot(\widetilde{J}n_{0})=0$ on $\Omega_{0}$.
As mentioned previously, Gauss/Appell is instantaneous. The decision variable is $a=\ddot{X}$; $X$ and $\dot{X}$ are frozen data. Thus feasibility (position/velocity level) is assumed, not solved for, at $t_{0}$. Differentiating the feasibility relations in time and evaluate at $t_{0}$:
$\widetilde{J}_{ij}\,\partial_{x_{0}j}a_{i}\;=\;-\,\dot{\widetilde{J}}_{ij}\,\partial_{x_{0}j}\dot{X}_{i}\qquad\text{in }\Omega_{0},$ (13)
$a\!\cdot(\widetilde{J}n_{0})\;+\;\dot{X}\!\cdot(\dot{\widetilde{J}}\,n_{0})\;=\;0\qquad\text{on }\partial\Omega_{0},$ (14)
which reduces to $a\!\cdot(\widetilde{J}n_{0})=0$ in the far-field when $\dot{\widetilde{J}}\to 0$.
Least–constraint
Given an impressed material acceleration $b(\xi,\eta)$, define
$\mathcal{G}[a]\;=\;\frac{1}{2}\int_{\Omega_{0}}\rho\,|a-b|^{2}\,dA,\qquad\text{with }X,\ \dot{X}\ \text{frozen at }t_{0}.$ (15)
We introduce the interior multiplier $\lambda(\xi,\eta)$ for (13) and boundary multipliers $\mu$ on $\partial\Omega_{0}$. (We do not add multipliers for $J=1$, $\dot{J}=0$, $u\!\cdot n=0$, since those are feasibility of $(X,\dot{X})$.)
$\mathcal{L}[a,\lambda,\mu_{w},\mu_{\infty}]=\int_{\Omega_{0}}\Big\{\tfrac{1}{2}\,\rho\,|a-b|^{2}+\lambda\big(\widetilde{J}_{ij}\partial_{x_{0}j}a_{i}+\dot{\widetilde{J}}_{ij}\partial_{x_{0}j}\dot{X}_{i}\big)\Big\}\,dA+\int_{\partial\Omega_{0}}\mu\big(a\!\cdot(\widetilde{J}n_{0})+\dot{X}\!\cdot(\dot{\widetilde{J}}n_{0})\big)\,ds_{0}.$ (16)
Write the interior term explicitly:
$\widetilde{J}_{ij}\,\partial_{x_{0}j}a_{i}=Y_{\eta}\,a_{x,\xi}-Y_{\xi}\,a_{x,\eta}-X_{\eta}\,a_{y,\xi}+X_{\xi}\,a_{y,\eta}.$
Applying the first variation and spatial integration by parts in $\xi$ and $\eta$ gives the interior Euler–Lagrange equations
${\begin{aligned} \rho(a_{x}-b_{x})-(\lambda Y_{\eta})_{\xi}+(\lambda Y_{\xi})_{\eta}&=0,\\ \rho(a_{y}-b_{y})+(\lambda X_{\eta})_{\xi}-(\lambda X_{\xi})_{\eta}&=0,\end{aligned}}\qquad\text{in }\Omega_{0},$ (17)
and boundary compatibility $\quad\lambda+\mu=0\quad\text{on }\partial\Omega_{0}\quad$}.
$\delta\lambda\ \ :\ \ {\quad Y_{\eta}a_{x,\xi}-Y_{\xi}a_{x,\eta}-X_{\eta}a_{y,\xi}+X_{\xi}a_{y,\eta}=-\,\dot{\widetilde{J}}_{ij}\partial_{x_{0}j}\dot{X}_{i}\quad\text{in }\Omega_{0}.\quad}$ (18)
$\delta\mu:\ \ {\quad a\!\cdot(\widetilde{J}n_{0})+\dot{X}\!\cdot(\dot{\widetilde{J}}n_{0})=0\quad\text{on }\partial\Omega_{0}\quad}$ (19)
Thus we have, for feasibility at $t_{0}:$ $J=1,\qquad\widetilde{J}_{ij}\partial_{x_{0}j}\dot{X}_{i}=0\quad\text{in }\Omega_{0},\qquad\dot{X}\!\cdot(\widetilde{J}n_{0})=0\quad\text{on }\partial\Omega_{0}.$ Therefore,
(acceleration level by Gauss)
$\widetilde{J}_{ij}\partial_{x_{0}j}a_{i}$ $=-\,\dot{\widetilde{J}}_{ij}\partial_{x_{0}j}\dot{X}_{i}$ $\text{in }\Omega_{0},$ (20)
$a\!\cdot(\widetilde{J}n_{0})+\dot{X}\!\cdot(\dot{\widetilde{J}}n_{0})$ $=0$ $\text{on }\partial\Omega_{0}.$ (21)
[…]
$\lambda+\mu$ $=0$ $\text{on }\partial\Omega_{0}.$ (23)
Transforming to Eulerian Coordinates: It is more intuitive to write these expressions in the Eulerian form: Let $x=X(\xi,\eta,t_{0})$, $J=1$, $u=\dot{X}^{\sharp}$, $a=\ddot{X}^{\sharp}=D_{t}u$, $P_{R}=\lambda^{\sharp}$, and $b^{\sharp}=b\circ X^{-1}$. We will use Piola/Nanson transformations (Gurtin, 2010):
$\operatorname{Div}_{0}(\lambda\,\widetilde{J})\mapsto\nabla_{x}P_{R},\qquad\widetilde{J}_{ij}\partial_{x_{0}j}(\cdot)\mapsto\nabla_{x}\!\cdot(\cdot)^{\sharp},\qquad(\widetilde{J}n_{0})\,ds_{0}\mapsto n\,ds,$
where the push forward transformation $\mapsto$ and superscript $(\cdot)^{\sharp}$ denote the transfer from reference coordinates to current coordinates.
(feasible at $t_{0}$)
$\nabla_{x}\!\cdot u=0\ \text{ in }\Omega,\qquad u\!\cdot n=0\ \text{ on }\partial\Omega_{0}.$ (24)
(acceleration level)
$\nabla_{x}\!\cdot(D_{t}u)$ $=\nabla_{x}\!\cdot\big((u\!\cdot\!\nabla_{x})u\big)$ $\text{in }\Omega,$ (25)
$D_{t}u\!\cdot n$ $=(u\!\cdot\!\nabla_{x}u)\!\cdot n$ $\text{on }\partial\Omega_{0}.$ (26)
$\rho\,(D_{t}u-b^{\sharp})+\nabla_{x}P_{R}$ $=0$ $\text{in }\Omega,.$ (27)
If $b^{\sharp}=-(1/\rho)\nabla_{x}P_{F}$, then $p=P_{F}+P_{R}$ and $\rho(\partial_{t}u+(u\!\cdot\!\nabla)u)=-\nabla_{x}p$.
Given (Lagrangian wall constraint at the instant $t_{0}$)
$a\cdot(\widetilde{J}\,n_{0})\;+\;\dot{X}\cdot\big(\dot{\widetilde{J}}\,n_{0}\big)\;=\;0.$ (28)
We use two identities Gurtin (2010) in the push–forward. At $t_{0}$:
$\text{(Nanson)}\qquad\widetilde{J}\,n_{0}\,dS_{0}\;=\;J\,n\,dS,$ (29)
$\text{(cofactor time derivative)}\qquad\dot{\widetilde{J}}\;=\;-\,(\nabla_{x}u)^{\!T}\,\widetilde{J}$. (30)
Multiply (28) by the reference line element $dS_{0}$ and use (29)–(30):
$\big[a\cdot(\widetilde{J}n_{0})\big]\,dS_{0}\;=\;a\cdot n\,dS,$
$\big[\dot{X}\cdot(\dot{\widetilde{J}}n_{0})\big]\,dS_{0}\;=\;u\cdot\big(-(\nabla_{x}u)^{\!T}\widetilde{J}n_{0}\big)\,dS_{0}\;=\;-\,u\cdot\big((\nabla_{x}u)^{\!T}n\big)\,dS,$
$\textrm{Hence }\big(a\cdot n-u\cdot[(\nabla_{x}u)^{\!T}n]\big)\,dS=0,\qquad\Rightarrow\qquad a\cdot n\;=\;u\cdot\big[(\nabla_{x}u)^{\!T}n\big]\ \ \text{on the wall.}\,$
Because $a=\partial_{t}u+(u\!\cdot\!\nabla_{x})u$ and $(\nabla_{x}u)\,u=(u\!\cdot\!\nabla_{x})u$,
$u\cdot[(\nabla_{x}u)^{\!T}n]\;=\;\big((\nabla_{x}u)\,u\big)\cdot n\;=\;(u\!\cdot\!\nabla_{x}u)\cdot n,$
so $(\ast)$ is equivalent to
$(\partial_{t}u+(u\!\cdot\!\nabla_{x})u)\cdot n\;=\;(u\!\cdot\!\nabla_{x}u)\cdot n\rightarrow u_{t}\cdot n=0.$