### v2

# Gauss Principle in Incompressible Flow: A Unified Variational Perspective on Pressure and Projection

Abstract
Following recent work (Gonzalez & Taha, 2022; Peters & Ormiston, 2025), this manuscript clarifies what the Gauss–Appell principle determines in incompressible, inviscid flow and how it connects to classical projection methods. At a fixed time, freezing the velocity and varying only the material acceleration leads to minimization of a quadratic subject to acceleration-level constraints. First-order conditions yield a Poisson–Neumann problem for a reaction pressure whose gradient removes the non-solenoidal and wall-normal content of the provisional residual, i.e. the Leray–Hodge projection. Thus, Gauss–Appell enforces the instantaneous kinematic constraints and recovers Euler at the instant. In steady flows, this principle –on its own– cannot select circulation or stagnation points because these are properties of the velocity state, not the instantaneous acceleration correction. The principle only determines the reaction pressure for an already-specified velocity field. The impressed/reaction pressure bookkeeping can be supplemented with orthogonality conventions that separate prescribed conservative forcing (if any) from the reaction enforcing the constraints. This variational viewpoint also yields a simple computational diagnostic: the minimized Appellian equals a $L^{2}$ norm of the reaction-pressure gradient which vanishes for constraint-compatible updates and grows with the magnitude of divergence and wall-flux mismatch. The goal of this note is simply to lend more clarity to the application of the Gauss principle, and to connect it concretely to well known concepts including potential flow theory, recent variational approaches and projection algorithms.
keywords:
Variational theory; Principle of least constraint; Gauss principle
The application of variational principles to fluid mechanics has a rich history spanning over 150 years, beginning with Helmholtz’s vortex theorems (1858) and Kelvin’s circulation theorem (1869). Classical approaches typically employed Hamilton’s principle, extremizing action integrals over time to derive the equations of motion through calculus of variations (Holm et al., 1998; Truesdell & Noll, 2004; Marchioro & Pulvirenti, 2012). These time-integral formulations, while mathematically elegant, vary entire space-time paths and thus provide limited insight into the instantaneous mechanics of constraint enforcement in incompressible flows  (Holm et al., 1998; Truesdell & Noll, 2004; Marchioro & Pulvirenti, 2012).
A fundamentally different variational approach emerged from Gauss’s principle of least constraint (Gauss, 1829), originally developed for constrained mechanical systems. Unlike Hamilton’s principle which varies paths over time, Gauss’s principle operates at a fixed instant, minimizing a quadratic form in accelerations subject to acceleration-level constraints. This instantaneous character makes it particularly suited to understanding how pressure maintains incompressibility moment by moment. Despite its potential, the principle remained relatively uncommon in classical fluid mechanics for over a century, even though it is widely used in constrained molecular dynamics (e.g. Nosé (1984); Hoover (1985); Evans & Morris (1990); Thijssen (2007). Recent works (Gonzalez & Taha, 2022; Taha et al., 2023; Taha & Shorbagy, 2025; Peters & Ormiston, 2025) have brought back a variational lens on fluid flows, particularly revisiting the Gauss principle in the context of the airfoil lift problem.
Despite the renewed interest, there remains confusion about what a Gauss/Appell least‑constraint statement actually determines in fluids, and how it can be utilized. We provide a compact variational interpretation of the pressure Poisson equation, clarify how to embed physical inputs via an ‘impressed’ subspace, and give a corollary that places the projection method in numerical hydrodynamics (Chorin, 1968) under Gauss/Appell.
Gonzalez & Taha (2022) minimize a steady reduced functional on a kinematically admissible manifold (e.g., a potential‑flow family parameterized by circulation), a model‑reduction approach but not the instantaneous Gauss principle. Taha et al. (2023) derive a general “minimum pressure-gradient” (PMPG) interpretation directly from Gauss’s principle (for arbitrary viscous stresses), showing that Navier–Stokes arises as the first-order optimality condition when the variation is taken with respect to the local acceleration. Steady/inviscid specializations of such costs have also been explored as selection mechanisms layered on top of Euler (e.g. Taha & Shorbagy (2025)). In particular, the fixed-time Gauss/Appell projection step considered here does not, by itself, resolve classical indeterminacies of steady exterior potential flow, because those indeterminacies live in the choice of steady velocity state rather than in the instantaneous constrained acceleration correction (see also Peters & Ormiston (2025)).
This work clarifies these viewpoints by showing precisely where Gauss/Appell ends (fixed‑time projection) and where additional modeling/selection begins (steady‑state choice of manifold or boundary data). The Gauss principle is a fixed-time variational statement: one freezes the velocity field and varies the material acceleration. Minimizing a positive‑definite “Appellian” in the admissible accelerations yields first order optimality conditions that are exactly the Leray–Hodge projection of the residual onto gradient fields; the Lagrange multiplier is a reaction pressure obtained from a Poisson–Neumann problem whose gradient removes the non‑solenoidal part of the residual. In steady flow, this principle does not select among kinematically admissible steady states; it only determines the pressure compatible with the given velocity field. These points are well known from projection methods but are often obscured in recent discussions; we make the equivalence and its scope precise.
Further, we clarify that shifting a known scalar potential between the two terms in $p=P_{F}+P_{R}$ is a gauge freedom in bookkeeping and does not imply physical non-uniqueness of the reaction (constraint) pressure. Building on Peters & Ormiston (2025), we show how an energy (Dirichlet) orthogonality condition can be used as a gauge-fixing convention to pick a unique representative once a reference/impressed subspace is chosen, and connect this to the classical Leray projection. When $P_{F}$ is reserved for truly impressed physics (e.g. conservative body-force potentials specified independently of incompressibility), the split is unique and $P_{R}$ is the unique Lagrange multiplier enforcing the constraints (up to a constant).
Unlike the principle of least action, Gauss/Appell is a least‑deviation principle in the accelerations at a given instant. In this sense, it is clarified that Gauss/Appell provides a unifying variational interpretation of the pressure Poisson solve—useful analytically and practically, particularly when one wishes to embed physical inputs (e.g., circulation or far‑field content) through an impressed subspace while keeping the constraints exact. We present rigorous derivations and provide insight into the physical implications of the Gauss principle.
1 Setting, notations
Let $\Omega$ be the exterior domain of a smooth body, and $\partial\Omega$ be the boundaries. The density $\rho$ is constant. $u(\cdot,t)$ is the velocity field. The Gauss’ principle is applied at a fixed time. One freezes the state $u(\cdot,t)$ and chooses an admissible instantaneous update in acceleration. For fluids, the material acceleration is $a\triangleq\partial_{t}u+(u\!\cdot\!\nabla)u.$ We assume throughout that $u(\cdot,t)$ is sufficiently smooth such that the quadratic functional below is finite. At a fixed time, the convective part $(u\!\cdot\!\nabla)u$ is a known function of space (because $u(\cdot,t)$ is frozen). Hence a perturbation of acceleration comes entirely from perturbing $u_{t}$: $\delta a=\delta u_{t}$. This is the fluid analogue of Gauss in constrained multibody dynamics (Udwadia & Kalaba, 1992), where one minimizes a positive-definite quadratic form in accelerations at a fixed configuration and velocity, subject to acceleration-level constraints; here the Eulerian fixed-time data are $(u,\nabla\!\cdot u,u\!\cdot n)$. Therefore, the variational variable is $u_{t}(\cdot,t)$, not a time-slab history, and not a configuration displacement as in stationary action.
For an airfoil, with far-field velocity $U_{\infty}$ and slip wall:
$\nabla\!\cdot u=0\quad\text{in }\Omega,\qquad u\!\cdot n=0\quad\text{on }\partial\Omega,\qquad u\to U_{\infty}\text{ as }|x|\to\infty.$
At the fixed instant $t$, these are given. Differentiating the two kinematic constraints in time gives the acceleration-level constraints on $u_{t}$:
$\nabla\!\cdot u_{t}=0\quad\text{in }\Omega,\qquad u_{t}\!\cdot n=0\quad\text{on }\partial\Omega.$ (1)
Below, the Eulerian form of the derivation is detailed.
1.1 Impressed vs reaction pressure and the Appellian
Motivated by Peters & Ormiston (2025), we write the total pressure
$p\triangleq P_{F}+P_{R},$
where
  • $P_{F}$ (impressed / prescribed) is a scalar potential associated with conservative impressed forcing that is specified independently of the incompressibility and wall constraints (e.g. a body-force potential); in unforced Euler one may take $P_{F}\equiv 0$.
  • $P_{R}$ (constraint) is the Lagrange–multiplier that enforces the acceleration–level constraints completely and uniquely. Its gradient $\nabla P_{R}$ is the constraint force per unit volume.
For a fixed velocity snapshot $u(\cdot,t)$ and prescribed impressed potential $P_{F}$, the reaction pressure $P_{R}$ is determined uniquely (up to an additive constant) by the Poisson–Neumann problem (5). In the linearized unforced setting emphasized in Peters & Ormiston (2025), taking the divergence yields $\nabla^{2}p=0$, so only the sum $p=P_{F}+P_{R}$ is constrained to be harmonic. The stronger separation $\nabla^{2}P_{F}=\nabla^{2}P_{R}=0$ (their Eq. (17)) is therefore an additional harmonic-gauge convention rather than a restriction on physically impressed forces, which are prescribed independently of the incompressibility constraint. In the present work we do not infer such restrictions on $P_{F}$; once any impressed conservative forcing is specified, the reaction pressure $P_{R}$ enforcing (1) is uniquely determined (up to a constant) by (5).
Define the frozen field
$C\triangleq(u\!\cdot\!\nabla)u+\frac{1}{\rho}\,\nabla P_{F},$ (2)
which is entirely known once $u(\cdot,t)$ and $P_{F}(\cdot,t)$ are specified. Gauss/Appell at the instant minimizes the Appellian
[…]
The quadratic form in (3) uses the mass-weighted $L^{2}$ inner product on accelerations, and yields the classical Leray–Hodge projection. In the function-space setting this choice is not unique: other metrics define different least-constraint projections. Weaker norms and lower regularity settings have been used in variational approaches to high-Reynolds-number flows; see e.g. Hoffman & Johnson (2010); Hoffman et al. (2016).
2 First order stationarity
We write the constrained Lagrangian and derive the first order Karush-Kuhn-Tucker (KKT) conditions. The Lagrangian with multipliers $\lambda$ (volume) and $\mu$ (boundary) is:
$\mathcal{L}[u_{t},\lambda,\mu]\triangleq\frac{1}{2}\int_{\Omega}\rho\,|u_{t}+C|^{2}\,dV\;-\;\int_{\Omega}\lambda\,(\nabla\!\cdot u_{t})\,dV\;-\;\int_{\partial\Omega}\mu\,(u_{t}\!\cdot n)\,dS.$
Variations yield
$\delta_{\lambda}:\ \nabla\!\cdot u_{t}=0,\qquad\delta_{\mu}:\ u_{t}\!\cdot n=0,$
and
$\delta_{u_{t}}\mathcal{L}$ $=\int_{\Omega}\big(\rho(u_{t}+C)+\nabla\lambda\big)\cdot\delta u_{t}\,dV\;+\;\int_{\partial\Omega}(\lambda-\mu)\,\delta u_{t}\!\cdot n\,dS.$
Because $\delta u_{t}$ is arbitrary in the interior and its normal component is arbitrary on $\partial\Omega$ (before enforcing the constraint), stationarity gives
$\rho(u_{t}+C)+\nabla\lambda=0\quad\text{in }\Omega,\qquad\mu=\lambda\quad\text{on }\partial\Omega,$ (4)
together with (1). This derivation follows the same constrained-optimization/projection-step logic as Peters & Ormiston (2025), while emphasizing that in the present work $P_{F}$ denotes only prescribed conservative impressed forcing (if any) and the reaction pressure $P_{R}$ carries all constraint-enforcing pressure content.
Now, define $P_{R}\triangleq\lambda$. From (4), $u_{t}=-(C+\nabla P_{R}/\rho)$; impose (1):
$\nabla\!\cdot u_{t}=0\ \Rightarrow\ \nabla^{2}P_{R}=-\,\rho\,\nabla\!\cdot C\quad\text{in }\Omega,$
and
$u_{t}\!\cdot n=0\ \Rightarrow\ \partial_{n}P_{R}=-\,\rho\,C\!\cdot n\quad\text{on }\partial\Omega.$
Thus the first step is the scalar Poisson–Neumann boundary value problem
$\boxed{\;\nabla^{2}P_{R}=-\,\rho\,\nabla\!\cdot C\ \text{ in }\Omega,\qquad\partial_{n}P_{R}=-\,\rho\,C\!\cdot n\ \text{ on }\partial\Omega}$ (5)
Compatibility holds identically: $\int_{\Omega}\nabla^{2}P_{R}\,dV=\int_{\partial\Omega}\partial_{n}P_{R}\,dS$ because $\int_{\Omega}\nabla\!\cdot C\,dV=\int_{\partial\Omega}C\!\cdot n\,dS.$ With (5) solved, we have $u_{t}^{\star}=-(C+\nabla P_{R}/\rho),\qquad p=P_{F}+P_{R}.$ Since $a=u_{t}+(u\!\cdot\!\nabla)u$, (4) is equivalent to
$\rho a=-\nabla p,$
i.e. Euler at the instant.
Once optimized, the minimized Appellian is purely “projection effort”
$\mathcal{A}_{\star}=\frac{1}{2}\int_{\Omega}\frac{1}{\rho}\,|\nabla P_{R}|^{2}\,dV.$
This reduced functional is the instantaneous $L^{2}$ norm of the reaction pressure gradient (the “minimum pressure-gradient” quantity emphasized by Taha et al. (2023)). This is precisely the standard projection method logic in continuous form: solve a Poisson–Neumann problem for pressure and project the residual onto gradients. It is notable that the Leray projector (see, for instance, Temam (2024)) maps any field to the divergence-free, wall-tangent subspace.
Writing $R\triangleq\rho\ a+\nabla P_{F}=\rho(u_{t}+C),$ the KKT conditions $\rho(u_{t}^{\star}+C)=-\nabla P_{R}$ says $R=-\nabla P_{R}$: the solenoidal, wall-tangent part of $R$ vanishes. Thus $P_{R}$ realizes the Leray projection of $R$ onto gradients, and 5 is the strong-form projector.
3 Physical implications and insight
At a fixed instant, the Gauss–Appell step has a concrete physical meaning: it chooses the smallest admissible change to the material acceleration that restores the instantaneous momentum balance while preserving incompressibility and wall non-penetration. The only agency it has is a reaction field, $\nabla P_{R}/\rho$, whose role is to remove whatever divergence and normal flux the provisional residual carries. The pressure is a reaction that instantaneously enforces the kinematics; it does no work on divergence-free, wall-tangent motions. This point is correctly argued by  Gonzalez & Taha (2022).
3.1 Projection effort as a diagnostic
After elimination, the minimized Appellian $\mathcal{A}^{\ast}$ becomes a quantitative measure of the projection effort, i.e. how hard the constraint must “push” to keep the instantaneous update compatible with incompressibility and wall tangency. The minimized Appellian can be written as
$\mathcal{A}_{\star}=\frac{1}{2\rho}\,\|\nabla P_{R}\|_{L^{2}(\Omega)}^{2},$
and therefore provides a scalar measure of the magnitude of the constraint correction at a given instant. In particular, if the frozen residual $C$ is already compatible with the acceleration constraints, $\nabla\!\cdot C=0$ in $\Omega$ and $C\!\cdot n=0$ on $\partial\Omega$, then (5) reduces to Laplace’s equation with homogeneous Neumann data and hence $\nabla P_{R}\equiv 0$ (up to an additive constant), so that $\mathcal{A}_{\star}=0$. In computations, a sudden growth of $\mathcal{A}_{\star}$ indicates that the provisional update violates divergence or wall-flux constraints and the projection is applying a large correction; this may occur, for example, due to inconsistent boundary treatment or under-resolution. More details on the projection operator in the context of general boundary conditions can be found in Maria Denaro (2003).
3.2 Steady flow
It is emphasized that in a steady flow, the Gauss principle of least constraint finds the gradient one must subtract from the “residual” $\rho(u\cdot\nabla)u+\nabla P_{F}$ so that the momentum equation holds. There is no update to $u$ (the minimizer gives $u_{t}^{\ast}=0$) and no selection among multiple steady Euler fields that satisfy the kinematics. In other words, no selection of extra physics occurs.
Gonzalez & Taha (2022) restrict the velocity field to a steady potential-flow family $u(\cdot;\Gamma)\triangleq u_{0}+\Gamma u_{1}$ whose members satisfy continuity and wall tangency, and then define
$S(\Gamma)\triangleq\frac{1}{2}\int_{\Omega}\rho\,\big|\,u\!\cdot\!\nabla u\,\big|^{2}\,dV,$
seeking $\Gamma^{\star}=\arg\min_{\Gamma}S(\Gamma)$. This is a reduced-order optimization on a one-parameter manifold of kinematically admissible steady flows and appears to pick a $\Gamma$ consistent with the Kutta condition in nearly sharp airfoils. However, it does not directly follow from the instantaneous Gauss/Appell principle; it presupposes steadiness and a particular parameterization. Peters & Ormiston (2025) discuss cases where such criteria can select steady states that do not align with physical reality.
In the full Gauss setting, the state $u$ is frozen, the variable is $u_{t}$, and the outcome is the constraint projection $P_{R}$2), not the selection of $\Gamma$. In the steady case, for every $\Gamma$ there exists a pressure that makes $(u(\cdot;\Gamma),p(\cdot;\Gamma))$ a steady Euler solution. Applying Gauss at the instant then yields $u_{t}^{\ast}=0$ for every $\Gamma$. Therefore, the direct fixed-time application of Gauss’s principle does not, by itself, favor a particular value of $\Gamma$; it simply returns the reaction pressure associated with each already-specified steady velocity state. In the steady case, the fixed-time Gauss step adds no additional closure beyond computing the compatible pressure (any selection over steady states requires additional modeling/criteria).
Taha et al. (2023) (see also Taha & Shorbagy (2025)) propose the ‘Principle of Minimum pressure gradient’, and offer reduced‑manifold selections that align with classical expectations from applying the Kutta condition on nearly sharp airfoils. An important distinction is required to interpret scope: Gauss/Appell provides the fixed‑time projection; selection over steady states requires additional modeling choices.
In the unsteady case, since $u_{t}$ (hence $a$) is varied with $u$ held fixed, acceleration-level constraints are enforced, resulting in the reaction pressure via the Poisson–Neumann problem (5). This exactly recovers Euler at the instant and clarifies the roles of impressed vs constraint pressure. This also explains why the fixed-time least-constraint step does not, by itself, resolve the classical indeterminacy of steady exterior potential flow: the indeterminacy lives in the state manifold, not in the instantaneous constrained acceleration update (cf. discussion in Peters & Ormiston (2025)). That indeterminacy lives in the state manifold, not in the instantaneous constrained acceleration update.
3.3 Freedom in the impressed/reaction pressure bookkeeping
Let $\psi$ be any prescribed smooth scalar field (a reference pressure / gauge potential). Define a new split $\widetilde{P}_{F}\triangleq P_{F}+\psi,\quad\widetilde{P}_{R}\triangleq P_{R}-\psi.$ Then
$\widetilde{C}=(u\!\cdot\!\nabla)u+\frac{1}{\rho}\nabla\widetilde{P}_{F}=C+\frac{1}{\rho}\nabla\psi,$
and the first order optimality condition becomes
$\rho\,(u_{t}^{\star}+\widetilde{C})+\nabla\widetilde{P}_{R}=\rho\,(u_{t}^{\star}+C)+\nabla P_{R}=0,$
so the optimizer $u_{t}^{\star}$ is invariant. Likewise $p=\widetilde{P}_{F}+\widetilde{P}_{R}$ is unchanged. Hence putting the circulation parameter $\Gamma$ “inside $P_{R}$” or “inside $P_{F}$” produces the same physical $(u_{t}^{\star},p)$; the split has a one‑parameter family of representatives. This is why $P_{R}$ is not unique a priori.
Similarly, the Poisson equation changes to $\nabla^{2}\widetilde{P}_{R}=-\rho\,\nabla\!\cdot C-\nabla^{2}\psi,\ \partial_{n}\widetilde{P}_{R}=-\rho\,C\!\cdot n-\partial_{n}\psi$. Therefore $\widetilde{P}_{R}=P_{R}-\psi+\text{const}$, exactly as above. So the Poisson step leaves $(u_{t}^{\star},p)$ invariant. This invariance reflects the classical fact that adding a prescribed gradient force (or, equivalently, redefining a reference pressure) does not affect the divergence-free, wall-tangent evolution; it is a gauge freedom in the representation of $p$, not a physical non-uniqueness of the reaction force. Once $P_{F}$ is fixed by the impressed physics (if any), the reaction pressure $P_{R}$ solving (5) is uniquely determined up to an additive constant. Once $P_{R}$ is solved, the optimal update is $u_{t}^{\star}=-(C+\nabla P_{R}/\rho)$, and the instantaneous Euler balance $\rho(\partial_{t}u+(u\!\cdot\!\nabla)u)=-\nabla(P_{F}+P_{R})$ is recovered.
3.4 Uniqueness of Dirichlet Orthogonality
Peters & Ormiston (2025) propose the following Dirichlet orthogonality as a convenient gauge-fixing convention for selecting a unique representative of the bookkeeping split $p=P_{F}+P_{R}$ (once a reference/impressed subspace has been chosen):
$\int_{\Omega}\nabla P_{R}\cdot\nabla P_{F}\,dV=0.$ (6)
It is stressed that (6) is not a physical restriction on an impressed force defined by external physics; it is a convention that fixes the gauge in the pressure bookkeeping. If $P_{F}$ is reserved for physically prescribed impressed potentials, the split is already unique (with $P_{R}$ the unique Lagrange multiplier enforcing the constraints) and (6) need not hold.
In Appendix A, it is shown that given any provisional solution $(\widehat{P}_{F},\widehat{P}_{R})$, the unique orthogonal pair is obtained by a correction:
$P_{R}=\widehat{P}_{R}-\sum_{j}\alpha_{j}\psi_{j},\qquad P_{F}=\widehat{P}_{F}+\sum_{j}\alpha_{j}\psi_{j}.$
The coefficients $\alpha_{j}$ are the orthogonal projection of $\widehat{P}_{R}$ onto the impressed subspace $\mathcal{S}_{\text{imp}}=\text{span}\{\psi_{1},\ldots,\psi_{m}\}$. We are subtracting this projection from $P_{R}$ (making it orthogonal to $\mathcal{S}_{\text{imp}}$) and adding it to $P_{F}$ (so the total pressure remains unchanged).
In other words, one can choose
  1. 1.
    What $\mathcal{S}_{\text{imp}}=\text{span}\{\psi_{1},\ldots,\psi_{m}\}$ is — which pressure modes one may want to impress (e.g., circulation patterns)
  2. 2.
    How much of each mode to impress, typically by specifying boundary conditions or forcing terms
The Dirichlet orthogonality condition then ensures consistent bookkeeping when decomposing the resulting total pressure $p$ (which emerges from the full flow physics). $\mathcal{S}_{\text{imp}}$ is the “signature” of the forcing mechanism. Dirichlet orthogonality thus labels as forcing ($P_{F}$) everything that has this signature; and labels as constraint ($P_{R}$) everything orthogonal to it (in the Dirichlet sense). We further note that the above is exactly analogous to the Helmholtz–Leray projection operator $\mathbb{P}$ operating on any given velocity field $w$:
$\mathbb{P}w\;=\;w-\nabla q,\qquad\nabla^{2}q=\nabla\!\cdot w\ \ \text{in }\Omega,\qquad\partial_{n}q=w\!\cdot\!n\ \ \text{on }\partial\Omega.$
Equivalently, $\mathbb{P}w$ is the unique vector that satisfies orthogonality $\int_{\Omega}\mathbb{P}w\cdot\nabla\phi\,\mathrm{d}x=0$. This is a well-known result (See for instance Sohr (2012)), but the connection to the impressed/constraint split is new. Appendix B clarifies another point relevant to the projection.
3.5 Physical impressed forces vs. reference (gauge) choices
For the exterior airfoil problem the distinction between physical impressed forces and bookkeeping/reference choices is immediate. The circulation degree of freedom in exterior potential flow resides in the harmonic subspace of the velocity (divergence-free, curl-free, wall-tangent in a multiply connected domain) and cannot be represented as $\nabla\psi$ of a single-valued scalar pressure potential; hence it cannot be embedded into $P_{F}$ as a physical pressure field. By contrast, conservative body-force potentials (e.g. gravity) may legitimately be placed in $P_{F}$ as prescribed impressed physics. If one introduces additional reference content in $P_{F}$ for diagnostics, then that choice is conventional.
  1. 1.
    Harmonic velocity modes (circulation) are not pressure gradients. In an exterior domain, circulation enters through the harmonic subspace of the velocity (divergence-free, curl-free, wall-tangent with nonzero loop integral). Such modes are not of the form $\nabla\psi$ for any single-valued scalar field, so they cannot be represented as a pressure potential in $P_{F}$ (nor in the reaction pressure $P_{R}$). In the fixed-time Gauss/Appell step, the circulation parameter $\Gamma$ appears only through the chosen frozen state $u(\cdot,t)$.
  2. 2.
    Conservative body-force potential. It is legitimate to include a prescribed uniform acceleration (e.g. gravity) through an impressed potential $P_{F}(x)=-\rho\,g\cdot x$, for which $(1/\rho)\nabla P_{F}=-g$ has units of acceleration. The Poisson step then returns $P_{R}$ to enforce $\nabla\cdot u_{t}=0$ and $u_{t}\cdot n=0$ for the frozen $u$.
  3. 3.
    Reference pressure for diagnostics (gauge). One may define a reference pressure for interpretation, e.g. subtract the Bernoulli pressure of a chosen reference potential flow in post-processing: $p^{\prime}:=p+\tfrac{1}{2}\rho|u_{\rm ref}|^{2}$. This is a change of variables (gauge) and should not be interpreted as an impressed force. For a given snapshot $u$ with no impressed forces, the pressure gradient is entirely the reaction that enforces incompressibility and wall compatibility (the Leray projection of $\rho(u\cdot\nabla)u$).
These remarks are particularly relevant to the discussion in Peters & Ormiston (2025) regarding “impressed pressure”. In the present framework, the frozen-time Gauss/Appell step returns
$u_{t}^{\star}=-\Big((u\!\cdot\!\nabla)u+\frac{1}{\rho}\nabla P_{F}+\frac{1}{\rho}\nabla P_{R}\Big),$
where $P_{F}$ is reserved for prescribed conservative impressed forcing (if any), and $P_{R}$ is the reaction (constraint) pressure. Enforcing the acceleration-level constraints $\nabla\!\cdot u_{t}^{\star}=0$ in $\Omega$ and $u_{t}^{\star}\!\cdot n=0$ on $\partial\Omega$ yields the Poisson–Neumann problem
$\nabla^{2}P_{R}=-\rho\,\nabla\!\cdot\Big((u\!\cdot\!\nabla)u+\frac{1}{\rho}\nabla P_{F}\Big),\qquad\partial_{n}P_{R}=-\rho\Big((u\!\cdot\!\nabla)u+\frac{1}{\rho}\nabla P_{F}\Big)\!\cdot n,$
which determines $P_{R}$ uniquely up to an additive constant. A genuinely impressed pressure potential corresponds to an externally prescribed conservative force $f=-\nabla\Phi$, for which one may take $P_{F}=\rho\Phi$; in the unforced exterior Euler airfoil setting $P_{F}\equiv 0$ and the entire pressure is reaction/constraint pressure.
Additionally, since $\nabla\times\nabla(P_{R}+P_{F})=0$, the pressure reaction is irrotational and thus does not appear explicitly in the interior vorticity evolution obtained by taking the curl of the momentum equation: $\omega_{t}+(u\!\cdot\!\nabla)\omega=(\omega\!\cdot\!\nabla)u$ (for conservative impressed forcing). We note that this statement concerns the bulk equation; boundary vorticity fluxes can be expressed in terms of tangential pressure gradients (e.g. Lighthill (1963) and Morton (1984).)
Finally, this viewpoint unifies interpretation and practice. The standard predict–Poisson–correct algorithm is precisely the Gauss–Appell step on a grid, and $\mathcal{A}^{\star}$ furnishes a diagnostic that can be monitored in computation: spikes may signal incompatible boundary data, under-resolved features, or an unhelpful choice of impressed/reference content. In reduced-order or data-assisted settings, the choice of $S_{\rm imp}$ allows trusted physics to be injected while the same projection guarantees exact enforcement of incompressibility and wall conditions.
Acknowledgments
Thanks to Robert Ormiston, Haithem Taha, Dave Peters and Philip Roe for discussions, and for inspiring this work.
Appendix A: Uniqueness of Dirichlet orthogonality
Substitute the correction 3.4 into the orthogonality condition 6:
$0=\int_{\Omega}\nabla P_{R}\cdot\nabla\psi_{i}\,dV=\int_{\Omega}\nabla\left(\widehat{P}_{R}-\sum_{j=1}^{m}\alpha_{j}\psi_{j}\right)\cdot\nabla\psi_{i}\,dV$ (7)
$\int_{\Omega}\nabla\widehat{P}_{R}\cdot\nabla\psi_{i}\,dV-\sum_{j=1}^{m}\alpha_{j}\int_{\Omega}\nabla\psi_{j}\cdot\nabla\psi_{i}\,dV$ (8)
Rearranging gives the $m\times m$ linear system:
$\sum_{j=1}^{m}\alpha_{j}\underbrace{\int_{\Omega}\nabla\psi_{j}\cdot\nabla\psi_{i}\,dV}_{G_{ij}}=\underbrace{\int_{\Omega}\nabla\widehat{P}_{R}\cdot\nabla\psi_{i}\,dV}_{b_{i}},\quad i=1,\ldots,m$ (9)
In matrix form: $G\bm{\alpha}=\mathbf{b}$, where:
  • $G_{ij}=\int_{\Omega}\nabla\psi_{j}\cdot\nabla\psi_{i}\,dV$ is the Gram matrix in the Dirichlet inner product $\langle p,q\rangle\triangleq\int_{\Omega}\nabla p\cdot\nabla q\,dV$.
  • $b_{i}=\int_{\Omega}\nabla\widehat{P}_{R}\cdot\nabla\psi_{i}\,dV$ projects the provisional reaction pressure onto each basis function
Once $\bm{\alpha}$ is computed by solving $G\bm{\alpha}=\mathbf{b}$, the unique Dirichlet-orthogonal decomposition is:
$P_{R}=\widehat{P}_{R}-\sum_{j=1}^{m}\alpha_{j}\psi_{j},\qquad P_{F}=\widehat{P}_{F}+\sum_{j=1}^{m}\alpha_{j}\psi_{j}$ (10)
Appendix B: Rank and Projection
Peters & Ormiston (2025) cast multipliers via a finite-dimensional Lagrange-multiplier theorem, tying the number of independent multipliers to the rank of a constraint Jacobian. In a discrete setting this perspective can be helpful. In the PDE setting, however, incompressibility and impermeability impose pointwise constraints. The Leray projector extracts the divergence-free, wall-tangent component, and the reaction pressure is recovered from the gradient component in the decomposition. The velocity space also contains a finite-dimensional harmonic subspace (e.g. circulation modes), which is left unchanged by the projector; see e.g. Sohr (2012) and the Euler/Appellian equivalence proof in Gonzalez & Taha (2022).
This viewpoint also makes the analogue of a “rank/consistency” condition explicit. The Neumann Poisson problem for $P_{R}$ is solvable if and only if the volume source and boundary flux satisfy the compatibility condition,
$\int_{\Omega}(-\,\nabla\!\cdot C)\,dV=\int_{\partial\Omega}(-\,C\!\cdot n)\,dS,$
which holds identically by the divergence theorem. Uniqueness is up to an arbitrary constant.
Physically, the decomposition expresses the familiar fact that pressure is a reaction field that enforces ideal kinematic constraints without doing work on admissible virtual velocities. Indeed, for any $v\in V$,
$\int_{\Omega}v\cdot\nabla P_{R}\,dV=\int_{\partial\Omega}P_{R}\,v\cdot n\,dS-\int_{\Omega}P_{R}\,\nabla\!\cdot v\,dV=0.$
Thus the reaction pressure is $L^{2}$-orthogonal to the space of admissible velocities, exactly as in d’Alembert’s principle, and the “multiplier” is most naturally interpreted as the gradient component required to project $\rho(u_{t}+C)$, rather than as a finite-dimensional rank count.